OpenStudy (anonymous):

find the locus of a point p(x,y) which moves so that its distance from the point (1,0) is half its distance from the line x=-1 whats my y point the second coordinate

OpenStudy (anonymous):

\[(x-1)^2 + (y)^2 = (x+1)^2/4(1)\]

OpenStudy (anonymous):

whats my y point? for d2?

OpenStudy (anonymous):

u dont have a y in the lines eqn so skip it

OpenStudy (anonymous):

y is arbritrary

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no..uve got a y on the left side havent u?

OpenStudy (anonymous):

so my final answer is x^2-2x+1+2y^2=0

OpenStudy (anonymous):

yes